Market demand is p=27x^(-1/3). The cost function of the monopolist is given by C(x)=3x. What is the MONOPOLIST's effective price (i.e. what is ps)? Keep fractions as long as possible to avoid rounding differences. Enter a number only, round to two decimals.
Answer #1

P = 27x-1/3 TR = Px = 27x+1/3.x = 27x2/3 MR = dTR/dx = 18x-1/3 TC = 3x MC = dTC/dx = 3 A monopolist produces output at the po